OpenStudy (anonymous):

find RANGE of function (2 x + 2)(x − 9) could u show me how to i forgot??

OpenStudy (anonymous):

the domain of the function is all valid inputs (x) the range of a function is all valid outputs (y) so another way to view this is what range of y values do you get from this?

OpenStudy (anonymous):

find the second coordinate of the vertex, that will be the minimum value you have to multiply out first and make it look like \(ax^2+bx+c\) then use \(-\frac{b}{2a}\) fo the first coordinate

OpenStudy (anonymous):

you should get \(2x^2-16x-18\) so vertex has first coordinate of \[-\frac{-16}{2\times 2}=4\] replace \(x\) by 4 to get the second coordinate of the vertex. since this is a parabola that opens up the range will be from there to \(\infty\)

OpenStudy (anonymous):

thanks satellite!